### 10 Mar 2026 v1

@@ 0.C.2 Scene Annotation Correction
For scene annotation correction, we first sample 5% of the data and ask annotators from three data-labeling companies to conduct pilot annotations. We then select the company with the highest annotation quality and commission it to annotate the full dataset. Annotators are required to cross-check each other’s work to ensure quality. During the annotation process, we hold multiple meetings with them to align on the annotation guidelines. After annotation is completed, our internal specialists review the quality and accept the deliverables only if they meet a 99% accuracy requirement. Below are the annotation guidelines provided to the human annotators from the data-labeling company. For brevity, we omit the examples.
@@ Appendix 0.D Detailed evaluation setups
LLM-based Scorers. To enable more researchers to run evaluations with our benchmark on a single consumer-grade GPU, we use an open-source LLM with no more than 10B parameters as the scorer. In addition to Qwen3-8B [yang2025qwen3], we also try other LLMs as the scorer, including Mistralai-3-8B-Instruct [mistral3] and Hunyuan-7B-Instruct [hunyuan]. Their scores achieve Spearman correlations ($\rho$) of 0.874 and 0.815 with human evaluations, respectively. As mentioned in the main text, using Qwen3-8B yields $\rho=0.919$, so we currently adopt Qwen3-8B as the scorer.
@@
Table 5: Evaluation results of Qwen3-VL-8B-Instruct with different inference strategies. Short, Medium and Long mean the subsets with short, medium and long atomic-action sequences, respectively. Full denotes the full dataset. Dark green and light green indicate the best and the second best result among all models.
@@ Appendix 0.E More Experimental Results
Quantitative Results of Inference Strategies. The quantitative results of Qwen3-VL-8B-Instruct with different inference strategies are reported in Tab.˜5.
Description Length. The visual comparisons of the average description length and unified score $S_{uni}$ on the full EXPLORE-Bench across different models are shown in Fig.˜8, which indicates that there is no clear correlation between description length and performance.
@@
Appendix 0.F More Cases
@@ 0.F.1 Cases of Final-scene Annotations
Here we present the annotations for the two final-scene images in Fig.˜9. LABEL:lst:anno_case1 contains the annotation of the final scene shown in Fig. 7 of the main text, so that readers can fully understand this case.
Listing 13: The scene annotation of case 1 (Fig.˜8(a))
@@
Listing 14: The scene annotation of case 2 (Fig.˜8(b))
@@
0.F.2 Case Study
Model-predicted final-scene descriptions are shown in Fig.˜10, Fig.˜11, and Fig.˜12.
Figure 10: Case study of case 1. “C” refers to the camera wearer. Note that the scene in this case is the same as in Fig. 7 of the main text, but we show predictions from different models, demonstrating their consistent failures in this case. Notably, Qwen3-VL-8B-Instruct and Keye-VL-1.5-8B cannot even realize that the bottle pyramid changes.
@@ 0.F.2 Case Study
Appendix 0.G Comparison with related works
In works like EgoSchema [mangalam2023egoschema], ERQA [team2025gemini], and EgoPlan-Bench[chen2026egoplan], actions are mainly represented through event descriptions or next-step prediction. Their action sequences are short and are primarily evaluated in a multiple-choice format (as shown in Fig.˜13). Embodied Agent Interface [li2024embodied] and EmbodiedBench [yang2025embodiedbench] evaluates MLLMs for embodied planning by focusing on end-to-end execution success. In contrast, our benchmark requires models to predict the causal effects of a sequence of atomic actions on the scene, with longer action sequences (as shown in Fig.˜10) and open-ended questions. Moreover, we provide a fine-grained, systematic scene-level evaluation across objects, attributes, and relations. Compared to ENACT’s [wang2025enact] symbolic forward/inverse world modeling (sequence reordering via scene-graph deltas in simulation), EXPLORE-Bench operates on real videos, conditions on long textual action sequences, and evaluates holistic visual end-state consistency, making it a complementary real-world testbed.
@@ Appendix 0.H Broader Impacts and Limitations
Broader Impacts. EXPLORE-Bench targets a capability central to embodied operation: anticipating the cumulative effects of actions. The measured deficits of current models, especially on abnormal and safety-critical cases, highlight gaps that must be closed before real-world deployment. We hope EXPLORE-Bench can catalyze progress in consequence-aware reasoning and safer egocentric agents.

### 12 Mar 2026 v2

@@ 0.C.2 Scene Annotation Correction
For scene annotation correction, we first sample 5% of the data and ask annotators from three data-labeling companies to conduct pilot annotations. We then select the company with the highest annotation quality and commission it to annotate the full dataset. Annotators are required to cross-check each other’s work to ensure quality. During the annotation process, we hold multiple meetings with them to align on the annotation guidelines. After annotation is completed, our internal specialists review the quality and accept the deliverables only if they meet a 99% accuracy requirement in spot-checking. Below are the annotation guidelines provided to the human annotators from the data-labeling company. For brevity, we omit the examples.
@@ Appendix 0.D Detailed evaluation setups
Human Evaluation Setup. Following VSI-Bench [yang2025thinking], during the evaluation of human performance on EXPLORE-Bench (tiny), human participants are allowed unlimited time to answer questions.
@@
LLM-based Scorers. To enable researchers to run evaluations on our benchmark using a single consumer-grade GPU, we use an open-source LLM with no more than 10B parameters as the scorer. In addition to Qwen3-8B [yang2025qwen3], we also try other LLMs as the scorer, including Mistralai-3-8B-Instruct [mistral3] and Hunyuan-7B-Instruct [hunyuan]. Their scores achieve Spearman correlations ($\rho$) of 0.874 and 0.815 with human evaluations, respectively. As mentioned in the main text, using Qwen3-8B yields $\rho=0.919$, so we currently adopt Qwen3-8B as the scorer.
@@
(a) Unified score vs. number of segments
(b) Unified score vs. window size
Figure 8: Unified score $\bm{S_{uni}}$ of Ovis2.5-9B across subsets under different inference strategies. Short, Medium, and Long denote the subsets with short, medium, and long atomic-action sequences. Full denotes the full dataset.
(a) Description length vs. number of segments
(b) Description length vs. window size
Figure 9: Average single-instance final-scene description length of Ovis2.5-9B under different inference strategies.
(a) Inference time vs. number of segments
(b) Inference time vs. window size
Figure 10: Average single-instance inference time of Ovis2.5-9B on an NVIDIA H200 GPU under different inference strategies.
Table 5: Evaluation results of Qwen3-VL-8B-Instruct under different inference strategies. Short, Medium and Long mean the subsets with short, medium and long atomic-action sequences, respectively. Full denotes the full dataset. Dark green and light green indicate the best and the second best result among all strategies.
@@
Table 6: Evaluation results of Ovis2.5-9B under different inference strategies. Short, Medium and Long mean the subsets with short, medium and long atomic-action sequences, respectively. Full denotes the full dataset. Dark green and light green indicate the best and the second best result among all strategies.
Partitioning Short Medium Long Full
Default Setting
[…]
Single-turn Inference with Various segment_num
segment_num = 11 34.30 0.75 1.63 26.87 32.43 0.77 1.67 26.88 30.10 0.75 1.69 26.31 33.14 0.76 1.66 26.81
segment_num = 10 34.16 0.74 1.63 26.81 33.33 0.80 1.70 27.55 31.76 0.80 1.76 27.64 33.59 0.77 1.67 27.18
segment_num = 9 33.89 0.74 1.63 26.71 33.75 0.81 1.74 27.98 30.69 0.78 1.75 27.11 33.48 0.77 1.68 27.23
segment_num = 8 33.91 0.77 1.66 27.12 34.22 0.82 1.74 28.19 31.84 0.81 1.78 27.89 33.80 0.79 1.70 27.60
segment_num = 7 35.28 0.82 1.73 28.39 35.09 0.85 1.79 29.05 33.25 0.84 1.84 28.87 34.98 0.83 1.76 28.69
segment_num = 6 36.24 0.87 1.80 29.56 36.54 0.91 1.86 30.39 34.55 0.90 1.91 30.18 36.16 0.89 1.84 29.94
segment_num = 5 37.83 0.90 1.85 30.56 37.47 0.92 1.91 31.11 34.86 0.90 1.92 30.34 37.37 0.91 1.88 30.74
segment_num = 4 39.51 0.96 1.95 32.25 39.79 1.02 2.00 33.07 36.85 0.96 1.99 31.91 39.32 0.98 1.98 32.52
segment_num = 3 41.59 1.05 2.05 34.12 42.05 1.07 2.11 34.88 38.66 1.01 2.07 33.27 41.43 1.05 2.07 34.31
segment_num = 2 44.37 1.15 2.16 36.43 43.02 1.13 2.14 35.75 39.06 1.07 2.11 34.14 43.28 1.13 2.15 35.92
Single-turn Inference with Various window_size
window_size = 10 35.96 0.86 1.82 29.57 30.74 0.72 1.61 25.55 29.45 0.70 1.66 25.60 33.30 0.79 1.72 27.64
window_size = 15 39.42 1.01 2.00 32.89 32.39 0.78 1.69 27.02 29.55 0.71 1.68 25.81 35.71 0.89 1.85 29.93
window_size = 20 41.09 1.04 2.05 33.99 33.86 0.83 1.77 28.43 31.06 0.76 1.73 26.98 37.30 0.93 1.91 31.15
window_size = 25 43.00 1.11 2.14 35.65 35.60 0.88 1.86 29.95 31.54 0.77 1.76 27.38 38.98 0.99 2.00 32.62
window_size = 30 43.86 1.15 2.15 36.24 36.72 0.92 1.90 30.83 32.74 0.82 1.84 28.70 39.98 1.03 2.02 33.40
window_size = 35 44.45 1.18 2.18 36.78 38.52 0.97 1.97 32.21 33.21 0.85 1.85 29.08 41.00 1.06 2.07 34.23
window_size = 40 44.70 1.20 2.20 37.23 39.33 0.99 2.00 32.79 33.51 0.83 1.86 29.08 41.47 1.08 2.09 34.68
window_size = 45 45.40 1.23 2.23 37.78 39.95 1.02 2.03 33.41 34.32 0.86 1.90 29.82 42.15 1.11 2.12 35.28
window_size = 50 45.45 1.24 2.24 38.02 41.01 1.05 2.07 34.16 34.27 0.85 1.96 30.23 42.56 1.13 2.15 35.72
window_size = 55 46.11 1.27 2.28 38.68 40.40 1.03 2.03 33.60 34.67 0.91 1.93 30.46 42.72 1.14 2.15 35.88
Mutil-turn Inference with Various segment_num
[…]
segment_num = 4 65.58 1.81 2.90 52.30 64.82 1.68 2.86 50.89 59.08 1.56 2.70 47.29 64.58 1.74 2.86 51.22
segment_num = 5 65.61 1.80 2.91 52.33 65.45 1.68 2.87 51.07 59.79 1.48 2.71 46.99 64.91 1.72 2.87 51.27
[…]
segment_num = 7 65.31 1.81 2.90 52.16 64.95 1.66 2.88 50.88 60.87 1.51 2.77 47.89 64.69 1.72 2.88 51.21
[…]
Mutil-turn Inference with Various window_size
window_size = 55 61.31 1.77 2.78 49.91 63.21 1.71 2.81 50.23 60.73 1.51 2.70 47.30 61.95 1.72 2.78 49.74
window_size = 50 61.52 1.77 2.78 50.02 63.21 1.71 2.83 50.37 60.90 1.54 2.78 48.27 62.08 1.72 2.80 49.96
[…]
window_size = 40 62.48 1.79 2.82 50.69 65.02 1.71 2.86 51.11 60.65 1.51 2.77 47.86 63.22 1.73 2.83 50.53
[…]
window_size = 20 64.58 1.79 2.87 51.65 64.96 1.67 2.87 50.91 60.85 1.53 2.77 48.09 64.31 1.72 2.86 50.98
window_size = 15 65.09 1.81 2.91 52.17 64.90 1.67 2.88 50.93 61.34 1.56 2.76 48.39 64.61 1.73 2.88 51.29
window_size = 10 65.02 1.80 2.91 52.10 64.71 1.71 2.89 51.30 60.53 1.54 2.76 47.98 64.41 1.74 2.88 51.34
@@ Appendix 0.E More Experimental Results
More Results of Inference Strategies. To strengthen the generality of our conclusions, we also evaluate Ovis2.5-9B [ovis2_5] (non-thinking mode) under two inference strategies. The trends shown in Fig.˜8, Fig.˜9, and Fig.˜10 are broadly consistent with those observed for Qwen3-VL-8B-Instruct [qwen3vl] (in main text Sec.˜5.2). Notably, under multi-turn inference, Ovis2.5-9B achieves a larger gain in $S_{uni}$ than Qwen3-VL-8B-Instruct: when seg_num=11, their $S_{uni}$ scores on the Full set increase by 4.22 and 1.69, respectively, relative to the default setting. However, the additional computational overhead remains non-negligible.
Quantitative Results of Inference Strategies. The quantitative results of Qwen3-VL-8B-Instruct and Ovis2.5-9B under different inference strategies are reported in Tab.˜5 and Tab.˜6, respectively.
Description Length. The visual comparisons of the average description length and unified score $S_{uni}$ on the full EXPLORE-Bench across different models are shown in Fig.˜11, which indicates that there is no clear correlation between description length and performance.
@@
Appendix 0.F More Cases and Case Studies
@@ 0.F.1 Cases of Final-scene Annotations
Here we present the annotations for the two final-scene images in Fig.˜12. LABEL:lst:anno_case1 contains the annotation of the final scene shown in Fig.˜7 of the main text, so that readers can fully understand this case.
Listing 13: The scene annotation of case 1 (Fig.˜11(a))
@@
Listing 14: The scene annotation of case 2 (Fig.˜11(b))
@@
0.F.2 Case Studies
Detailed case studies of model-predicted final-scene descriptions are shown in Fig.˜13, Fig.˜14, and Fig.˜15. A case study of stepwise reasoning implemented by single-turn inference vs. multi-turn inference is presented in Fig.˜16.
Figure 13: Case study of case 1. “C” refers to the camera wearer. Note that the scene in this case is the same as in Fig.˜7 of the main text, but we show predictions from different models, demonstrating their consistent failures in this case. Notably, Qwen3-VL-8B-Instruct and Keye-VL-1.5-8B cannot even realize that the bottle pyramid changes.
@@ 0.F.2 Case Studies
Figure 16: Case study of stepwise reasoning implemented by single-turn inference vs. multi-turn inference. In this case, segment_num=2. (1) and (2) denote the model-predicted descriptions of the intermediate scene and the final scene, respectively. The descriptions generated by multi-turn inference are clearly longer than those generated by single-turn inference, covering more objects and providing more detailed descriptions of object attributes and relations.
@@
Appendix 0.G Comparisons with related works
In works like EgoSchema [mangalam2023egoschema], ERQA [team2025gemini], and EgoPlan-Bench[chen2026egoplan], actions are mainly represented through event descriptions or next-step prediction. Their action sequences are short and are primarily evaluated in a multiple-choice format (as shown in Fig.˜17). Embodied Agent Interface [li2024embodied] and EmbodiedBench [yang2025embodiedbench] evaluates MLLMs for embodied planning by focusing on end-to-end execution success. In contrast, our benchmark requires models to predict the causal effects of a sequence of atomic actions on the scene, with longer action sequences (as shown in Fig.˜13) and open-ended questions. Moreover, we provide a fine-grained, systematic scene-level evaluation across objects, attributes, and relations. Compared to ENACT’s [wang2025enact] symbolic forward/inverse world modeling (sequence reordering via scene-graph deltas in simulation), EXPLORE-Bench operates on real videos, conditions on long textual action sequences, and evaluates holistic visual end-state consistency, making it a complementary real-world testbed.
@@ Appendix 0.H Broader Impacts and Limitations
Broader Impacts. EXPLORE-Bench targets a capability central to embodied operation: anticipating the cumulative effects of actions, which is aligned with practical needs in household agents. The measured deficits of current models, especially on abnormal and safety-critical cases, highlight gaps that must be closed before real-world deployment. We hope EXPLORE-Bench can catalyze progress in consequence-aware reasoning and safer egocentric agents.